OpenStudy (anonymous):

A child is creating a pyramid with building blocks. The top three levels include 5 blocks, 12 blocks, and 19 blocks. Part 1: How many blocks would be needed for a pyramid 25 levels tall? Part 2: Use complete sentences to explain how a sum of an arithmetic series was applied.

Parth (parthkohli):

We have an arithmetic sequence: 5,12,19,26.... where d = 7.

Parth (parthkohli):

Sum of an arithmetic sequence formula is applied, as the total number of blocks needed for \(n\) levels is \(5+12+19\cdots(\text{till }n \text{ terms})\)

OpenStudy (anonymous):

uhu...

Parth (parthkohli):

Did I confuse you? Sorry for that.

Parth (parthkohli):

Firstly, do you see how creating each level increases the number of blocks by 7?

OpenStudy (anonymous):

no im just agreeing with you

Parth (parthkohli):

Oh.

OpenStudy (anonymous):

n yes i understand so far

Parth (parthkohli):

So can you use the formula?\[Sum~of~an~arithmetic~sequence = { n \over 2} \left(a_1 + a_n \right) \]

Parth (parthkohli):

Here, the number of blocks(the number of terms) is 25. Substitute \(n\) with 25.

Parth (parthkohli):

Oh! And you have to find \(a_{25}\) before continuing. For that, you must use the \(n\)th term in an arithmetic sequence formula.\[a_{25}= a_1 + (25 - 1)d \]We know that a1 = 5, d = 7, n = 25.\[a_{25}= 5 + 24\cdot 7 \]

OpenStudy (anonymous):

so a25=203

Parth (parthkohli):

Yes. Now just find\[{25 \over 2}(5 + 203)\]

OpenStudy (anonymous):

which is 2600

Parth (parthkohli):

And that's it.

OpenStudy (anonymous):

thnxz ur a genius!